THEORY OF r_{m}^{(2)} STRUCTURAL PARAMETERS

\underline{JAMES K. G. WATSON}, Steacie Institute for Molecular Sciences, National Research Council, Ottawa, Ontario, Canada K1A OR6.

The r_{m}^{(2)} method described in the previous paper is discussed on the assumption that the isotopic changes in ϵ_{0} are adequately represented by the equation Δϵ_{0} = ∑_{i}(∂ϵ_{0}/∂m_{i})Δm_{i}. A formula is calculated for ∂ϵ_{0}/∂m_{i} in terms of the basic potential constants and coupling parameters of the molecule. Use of this formula then gives an approximate expression for the [r_{m}^{(2)} − r_{e}] differences in the structural parameters. Numerical tests of this expression are in progress.